### v1

# Don’t Use LLMs to Make Relevance Judgments

Don’t Use LLMs to Make Relevance Judgments
Ian Soboroff
National Institute of Standards and Technology [^1] This article reflects the views of the author and not necessarily those of NIST or of the U. S. Government
Gaithersburg
   Maryland    USA
ian.soboroff@nist.gov
(draft 2, Sep 9, 2024)
1 Introduction
Making the relevance judgments for a TREC-style test collection can be complex and expensive. A typical TREC track usually involves a team of six contractors working for 2-4 weeks. Those contractors need to be trained and monitored. Software has to be written to support recording relevance judgments correctly and efficiently.
The recent advent of large language models that produce astoundingly human-like flowing text output in response to a natural language prompt has inspired IR researchers to wonder how those models might be used in the relevance judgment collection process [2, 5].
At the ACM SIGIR 2024 conference, a workshop “LLM4Eval” provided a venue for this work, and featured a data challenge activity where participants reproduced TREC deep learning track judgments, as was done by Thomas et al [13]. I was asked to give a keynote at the workshop, and this paper presents that keynote in article form.
The bottom-line-up-front message is, don’t use LLMs to create relevance judgments for TREC-style evaluations.
2 Automatic evaluation
The idea of automatic evaluation for information retrieval came from a paper I wrote with Charles Nicholas and Patrick Cahan in 2001 [12]. I had been reading Ellen Voorhees well-known SIGIR paper from 1998 [14] which shows experimentally that people differ in their judgments of relevance, but those differences don’t affect the relative ordering of systems in a TREC evaluation. Surprised by this result, I wondered what would happen if the relevance judgments were randomly sampled from the pools. [^2] In TREC, the pools usually consist of the top $k$ unique documents from some or all of the runs submitted to the track. There are exceptions that incorporate manual searches or relevance feedback, but those innovations were not examined in the 2001 paper.
A representative example result from that work is shown in Figure 1. The single $+$’s are official scores from TREC, and the $\times$’s with whiskers up and down are scores obtained with random documents from the pool labeled as relevant. The points are ordered according to their official TREC scores. The key point to notice is that, using random judgments, the best systems look like the worst systems.
Figure 1: Sample result from [12], TREC-8, TREC-style pooling to depth 100.
Shortly thereafter, Jay Aslam and Robert Savell published a short paper [1] which explained my 2001 results. By selecting the relevant documents at random, I was implicitly selecting documents retrieved by many systems. So my 2001 paper shows the results of a popularity contest: the most frequently retrieved documents are considered to be relevant. Under this approach, the worst systems and the best systems both look bad, because they fail (or succeed) by retrieving documents that other systems don’t find.
Another way to think about Aslam and Savell is that by using the output of a system as the ground truth, I am measuring the similarity of the two systems, how close the retrieval system is to the model that created the ground truth.
These papers inspired a number of people to investigate the possibility of automatic evaluation. MacAvaney and Soldaini [8] is a recent instance that cites most of the work in this area. Also, around that time, the BLEU [9] automatic metric for machine translation and the ROUGE [7] automatic metric for summarization were published, and so automatic evaluation was of broad interest.
3 Uncertainty is the root of IR
Information retrieval is a field founded on inescapable uncertainty. We cannot fully model the meaning of the documents, because language is fundamentally ambiguous. We cannot fully comprehend the user’s notion of relevance, because that notion is defined with respect to the entire cognitive state of the person, which changes as they use the system. When we receive a query from a user, it is a poor representation of the user’s need or goal, and systems are forced to guess where to look. Industry systems make extensive use of behavioral observations, which are uncertain in their implications. So even though we have many tools, we cannot actually perform matching of documents to information needs with absolute certainty.
Retrieval models model that uncertainty. For example, BM25 [11] is a simplified term-weighting scheme inspired by a statistical model of relevance [10], specifically that terms in relevant and non-relevant documents have different distributions of the same shape. This is a model of the relationship of words in a document to relevance, and one that has been outstandingly successful in practice.
While early retrieval models developed prior to the web focused primarily on the content of the documents, web search engines quickly began to use data from the behavior of users of the search engine. Those behavioral signals are not related to the words in the document, but rather both the document and the behavior are clues about relevance, which is unknowable. By incorporating both sets of clues, these systems seem to outperform systems based on only one or the other.
For purposes of evaluation, in the Cranfield paradigm, we ask a human being whether the document is relevant to the information need, and we declare that to be the truth. Even though we know that relevance is fuzzy and time-sensitive and context-sensitive and idiosyncratic, we feel that asking a human about relevance is a close enough model of the ideal system. This person said these documents were relevant, and these other documents were not. If a system retrieved all of the documents that this human determined to be relevant ahead of those determined to not be relevant, then the system would obtain perfect performance for that person according to traditional metrics like precision and recall.
Pooling is a way to judge enough documents that in the end we have a human assessment that is better than any pooled system. Pooling works because the human sees documents from every system, subject to the pooling parameters. Any pooled system can at best have found all the relevant documents identified in the pool.
And so we evaluate the performance of a system moving within a fundamentally uncertain universe by taking part of that uncertainty, declaring it to be certainly true for purposes of evaluation, and moving on from there.
There are of course other approaches to overcoming this uncertainty. If we have many sets of relevance judgments, we could estimate a confidence interval on the systems performance subject to all of those “declarations” of relevance. Models that exploit behavioral cues are optimizing a model of relevance based on those cues, which vary from user to user, from session to session. We can imagine asking our human assessor about relevance in the context of their true task, at multiple points in their task progress, and develop a notion of relevance over time. But the uncertainty never entirely goes away.
4 Retrieval and evaluation are the same problem
A retrieval system, or a relevance model, is a model of relevance given available data. The system is trying to predict which documents are relevant and which documents are not. Even though real systems might try to optimize a pairwise or listwise output, or compute a degree of relevance of a document or a search engine result page, it is useful to think of all these processes as predicting relevance.
During relevance assessment, we are asking the assessor to decide whether documents are relevant or not. This, too, is essentially a prediction of relevance. It’s a well-informed prediction since the person is reading the document, but since the task is artificial, the assessor is basically saying that they would include this document in a report on the topic, a report which they don’t ever actually write. So we can call that a prediction too.
As with any prediction, there are errors of omission and commission (or Type I and Type II if you prefer), and those errors represent a maximum discriminative ability of those relevance judgments to distinguish systems.
We can use one set of predictions, the relevance judgments, to measure the performance of the other set of predictions, the system outputs. In doing so, we declare the relevance judgments to be truth. In fact, you can switch the two sets of predictions, declare the system output to be truth, and measure the “effectiveness” of the assessor compared to that of the system. All evaluations which compare a system output to an answer key are making a measurement with respect to the answer key, not with respect to the universe.
Since both retrieval systems and relevance assessors are making predictions of relevance, evaluation and retrieval are the same problem. We can imagine a very slow system that would have a human read every document and assess its relevance given a query. [^3] John Searle’s “Chinese Room” thought experiment (https://en.wikipedia.org/wiki/Chinese_room) posits a person in a box who receives questions through a slot and delivers answers out the slot. The questions and answers are in Chinese, a language which the person does not speak. Rather, the person follows a sophisticated set of instructions for generating an answer from a question, in Chinese, by manipulating symbols on the paper. Thus, the box appears to understand and communicate in human language, but is basically a computer. Having a person read the entire collection to make relevance judgments sounds like the worst Chinese Room ever. I could have a language model perform the same task: construct a prompt of a topic statement and a document, and ask for the LLM to say relevant or not, for every document in the collection. Asking the language model about relevance is the mirror of evaluation.
If we believed that a model was a good assessor of relevance, then we would just use it as the system. Why would we do otherwise? We don’t use human assessors that way, because it doesn’t scale. LLMs in 2024 don’t scale, but that feels like an engineering problem; we need better hardware and smaller models.
5 The ceiling on performance
The relevance judgments in a Cranfield experiment are a model of human and thus ideal performance. The goal of the system authors is to create a system that can understand relevance like a [specific] human does. If you take the relevance judgments and turn it into a run, it gets perfect scores on the appropriate metrics.
Historically, this was the goal of IR performance. IR systems are meant to augment people by scaling up their ability to understand information, and so the performance of people is the ideal.
This ideal is also a limit on what Cranfield can measure. If we imagine we have a system that is better than a human, for example by finding relevant documents that are not in the relevance judgments or correctly ranking a document which was assessed incorrectly, that system will score less than perfectly when we score it using the human relevance judgments. This has to be the case, because unjudged documents are assumed to not be relevant, and the documents found by this novel system are either absent from the relevance judgments or judged non-relevant when they should have been marked relevant. The system has retrieved documents which according to the evaluation relevance judgments are not relevant. And so this top-performing system is measured as performing less well than it does.
This is reflected in my 2001 paper, and with other papers that came later, but exemplified by Figure 1 above. The (true) top systems are under-ranked by the “model” of relevance. This must be true for any model of relevance that generates relevance judgments, be it human or machine. We cannot measure a system that is better than the relevance judgments. Or, rather, the evaluation can’t distinguish such a system from one that performs less than perfectly.
As a counter-example, in 2007 Büttcher et al trained a model using an incomplete set of manual judgments to classify a larger set of documents automatically, improving the collection [3]. In this case the evaluation model is privileged in comparison to the runs, in that it has relevance information that they do not. Relevance feedback nearly always improves performance, so we would expect a hybrid set of judgments like these to have the possibility of outperforming evaluation using the shallow judgments. This is outperforming the human, but doing so only with human relevance data.  [^4] We still have a grounding problem, in that you may not believe that the model makes accurate predictions. The process can be improved by doing a second stage of relevance assessments on the classifier outputs in order to estimate the classifier error rates.
And so when the relevance judgments are created by a person, the model can’t exceed the human ideal. If we had a model that had “super-human” performance, we would just make our IR system use that model. In the current state of the art, the most advanced LLMs are used as components of systems that may be hypothetically measured by relevance judgments generated from the same models. Those systems cannot perform better than the model generating the relevance judgments.
Obviously, the human that created the relevance judgments is not entirely ideal. The assessor is not all-knowing, all-seeing, all-reading with perfect clarity. Assessors make mistakes, and TREC participants are fond of finding them. More importantly, the assessor is only one person; someone else with the same information need would make different judgments. If we compared the assessor’s judgments to those of a secondary assessor by pretending that secondary assessor is a run, it would necessarily perform less than perfectly.
That means that even if we imagine that systems exist which perform better at the task than humans do, we can’t see that improved performance in a Cranfield-style evaluation. This may also be true for other kinds of evaluations, for example a user study where subjects hunt for the relevant documents, or an A/B test where the “correct” answers come from some key. It must be true for any evaluation where we are comparing a system output to a “gold standard,” for example in machine learning or natural language processing, because the gold standard represents ideal performance.
Lastly, the so-called “super-human” performance observed on benchmark datasets is actually just measurement error. Super-human performance would be scored as less than ideal by the established ground truth, because performing better than a human entails making different decisions than those in the ground truth.
To summarize, you should not create relevance judgments using a large language model, because:
  • You are declaring the model to represent ideal performance, and so you can’t measure anything that might perform better than that model.
  • The model used to create relevance judgments is certainly also used as part of the systems being measured. Those systems will evaluate as performing poorly even if they actually improve on the model, because improving on the model means retrieving new relevant unjudged documents.
  • When the next shiny model comes out, it will measure as performing less well than the old model, because it necessarily must retrieve unjudged documents or ones judged incorrectly as not relevant. And so the relevance judgments can only measure systems that perform worse than the state of the art at the time the relevance judgments were created. [^5] This might be a good thing. Since new models will look worse than old models, when their developers run them through leaderboard-style benchmarks, they will cast them aside because they seem to perform poorly. Then we won’t ever have to worry about having a new model.
6 A New Hope
Does that mean that LLMs are not useful for evaluation? No, it only means that using a model to construct the truth data is doomed to produce a poor evaluation. But we don’t have to do things that way.
If the evaluation model has access to privileged information, for example by being fine-tuned on manual relevance judgments on the evaluation topics, then those relevance judgments should still be able to measure systems that use the untuned model. This is what makes Büttcher et al’s [3] results work: the model is trained on relevance data, and so the model is more privileged than any system it is measuring. We actually already knew this: the fact that relevance feedback improves retrieval is a basic result in IR. If we have relevance information gleaned from many systems, essentially a fusion result, the outputs will perform better than any individual system and thus we can measure the individual systems.
This still has the problem that the evaluation isn’t future-proof: we might have a new model that outperforms the relevance feedback of the prior generation.
One simple idea that seems promising is to employ a LLM to follow the assessor and look out for mistakes. This can’t be done by simply asking, “Did you mean ‘relevant’ instead?” since people are primed to trust the computer more than they should. But it may be possible to automate a quality-control process using a model.
We might imagine using LLMs to automate subjects in a user study, or assist researchers coding their observations, or related ideas that might serve to scale up user studies. In these cases, the LLM is not generating the ground truth.
As long as we are not generating ground truth we will use to measure systems, we can use the LLMs to support evaluation. Generating a proxy for ground truth still fails: using relevance judgments to train a model of user satisfaction (or some other evaluation target not directly related to the relevance judgments), and using that model to measure the systems, for example.
7 Conclusions
I have discussed the limitations of using models to create relevance judgments. You don’t want to do that, because then you have limited what you can measure to the level of the generating model. If the generating model is also part of the evaluated systems, you are stuck in a loop, or perhaps falling into a bottomless pit.
This is similar to model collapse [6]. When you train the model using its own outputs, the performance of the model decreases. The collapse mechanism is the measurement error that comes from generating the truth data using the evaluated system. In this case, evaluation is a loss function computed based on the generated truth data.
This doesn’t mean that LLMs can’t allow us to do amazing things. As someone who got his start in IR working with LSI [4], which is essentially an optimal linear embedding, I am very excited by the idea of nonlinear embeddings. IR systems that use LLMs to surmount the vocabulary boundary have enormous promise for real users.
All models have limits, and humans do too. If we want to use the model to evaluate performance, we first need to consider if we are doing something past the ability of the model as used in that evaluation paradigm. The relevance judgments barrier is a fundamental limitation of evaluations that measure systems against ground truth.
8 Acknowledgments
I thank Ellen Voorhees and Rikiya Takehi for their comments on the talk and on early versions of this paper. I also thank the attendees of the SIGIR 2024 LLM4Eval workshop for their insightful questions.